OpenStudy (anonymous):

What is the slope of the line that passes through the points (–5, 3) and (7, 5)?

OpenStudy (anonymous):

You just have to use the equation \[y - y1 = m(x-x1)\]

OpenStudy (anonymous):

After finding the slope. do you know how to do that?

OpenStudy (anonymous):

not really I suck at math

OpenStudy (anonymous):

slope formula is \[m = \frac{ y_{2}- y_{1} }{ x_{2} - x _{1} }\]

OpenStudy (anonymous):

1/6

OpenStudy (anonymous):

thnx